# Transcendence and normality of complex numbers via Hurwitz continued fractions

### v2

@@ 3.1. HCFs Cylinders
Naturally, #$\#A$ stands for the number of elements in a set $A$. We denote by $\Omega(n)$ (resp. $\mathsf{R}(n)$, $\mathsf{Ir}(n)$, $\mathsf{EI}(n)$) the set of valid (resp. regular, irregular, extremely irregular) words of length $n$. We extend the notions regular, irregular, and extremely irregular to cylinders and prototype sets in an obvious way. It is well known that there are only finitely many prototype sets (see, for example, [16]). Actually, an open prototype set is one of the following sets or one of their right-angled rotations (see Figure 1):
@@
A significant part of our calculations and our arguments is based on an observation stemming from the definitions of cylinders: if $(a_{n})_{n\geq 1}$ is a sequence in $\mathscr{D}$, then
$\mathcal{C}_{1}(a_{1})=\mathfrak{F}\cap\iota\tau_{a_{1}}[\mathfrak{F}],\quad\mathcal{C}_{2}(a_{1},a_{2})=\mathfrak{F}\cap\iota\tau_{a_{1}}[\mathfrak{F}]\cap\iota\tau_{a_{2}}\iota\tau_{a_{1}}[\mathfrak{F}];$
@@
Since $\iota$ and $\tau_{a}$ are homeomorphisms from the Riemann sphere onto itself, we may replace $\mathfrak{F}$ and $\mathcal{C}$ with $\mathfrak{F}^{\circ}$ and $\mathcal{C}^{\circ}$, respectively, in (8). It follows from (8) that every factor of a regular word (finite or infinite) is again regular. We also deduce the next proposition.
@@ 3.2. Regular cylinders
We recall two important properties of regular cylinders. The first property tells us that every infinite sequence in $\mathscr{D}$ with large terms is regular. The second proposition establishes symmetries of the family of regular cylinders.
@@
and call $\mathsf{Di}_{8}$ the group of isometries they generate. Obviously, $\mathsf{Di}_{8}$ is the dihedral group of order $8$. Proposition 4.5 and Corollary 4.6 in [9] say that any $f\in\mathsf{Di}_{8}$ maps open cylinders onto open cylinders of the same level. These results are proven by explicitly determining the image of an open cylinder under $\mathsf{rot}$ and $\mathsf{mir}_{1}$. In particular, if $\mathsf{mir}_{2}:=\mathsf{rot}^{2}\mathsf{mir}_{1}$, $z\mapsto-\overline{z}$, and $\mathsf{Mir}\subseteq\mathsf{Di}_{8}$ is the subgroup generated by $\mathsf{mir}_{1}$ and $\mathsf{mir}_{2}$, we have the next result.
@@ 3.3. Irregular cylinders
The proof amounts to check that every unmentioned option for $\mathfrak{F}^{\circ}_{n}(\mathbf{a})\neq\varnothing$ satisfies
$\mathbf{a}b\in\Omega(n+1)\quad\text{ implies }\quad\mathbf{a}b\in\mathsf{R}(n+1),$
which follows from the stronger statement
@@
We rely on pictures in order to avoid long yet elementary computations.
@@
    When $\mathfrak{F}^{\circ}_{n}(\mathbf{a})=\mathfrak{F}^{\circ}\setminus\overline{\mathbb{D}}(1+i)$ we have $\iota[\overline{\mathfrak{F}}_{n}(\mathbf{a})]=\iota[\overline{\mathfrak{F}}]\setminus\mathbb{D}(1-i)$. We conclude from Figure 4 that $\iota\left[\overline{\mathfrak{F}}_{n}(\mathbf{a})\right]\cap\tau_{-b}\iota[\mathfrak{F}]\neq\varnothing$ if and only if $b\neq 1-i$, in which case $\iota\left[\mathfrak{F}^{\circ}_{n}(\mathbf{a})\right]\cap\tau_{-b}\iota[\mathfrak{F}^{\circ}]\neq\varnothing$. A similar argument holds when $\mathfrak{F}^{\circ}_{n}(\mathbf{a})=\mathfrak{F}^{\circ}\setminus i^{k}\overline{\mathbb{D}}(1+i)$ for some $k\in\{2,3,4\}$.
@@
    and, by Figure 5, $\iota\left[\overline{\mathfrak{F}}_{n}(\mathbf{a})\right]\neq\varnothing$ if and only if $\operatorname{Re}(b)\geq 0$. Figure 5 also shows that $\operatorname{Re}(b)\geq 0$ yields $\iota[\mathfrak{F}^{\circ}_{n}(\mathbf{a})]\cap\tau_{b}[\mathfrak{F}^{\circ}]\neq\varnothing$. A similar argument holds when $\mathfrak{F}^{\circ}_{n}(\mathbf{a})=\mathfrak{F}^{\circ}\setminus\overline{\mathbb{D}}(i)$.
@@
    When $\mathfrak{F}^{\circ}_{n}(\mathbf{a})=\mathfrak{F}^{\circ}\setminus(\overline{\mathbb{D}}(i)\cup\overline{\mathbb{D}}(-1))$, by Figure 6 we have $\tau_{b}\left[\mathfrak{F}^{\circ}_{n}(\mathbf{a})\right]\cap\iota[\overline{\mathfrak{F}}]\neq\varnothing$ if and only if $\operatorname{Re}(b)\geq 0$ and $\operatorname{Im}(b)\geq 0$. For such $b$ we have $\iota[\mathfrak{F}^{\circ}_{n}(\mathbf{a})]\cap\tau_{b}[\mathfrak{F}^{\circ}]\neq\varnothing$.
This shows that if $\mathbf{a}\in\mathsf{R}(n)$ and $\mathfrak{F}^{\circ}_{n}(\mathbf{a})\neq\varnothing$ is not one of the sets in the statement, then $\mathbf{a}b\in\Omega(n+1)$ implies $\mathbf{a}b\in\mathsf{R}(n+1)$. ∎
@@ 4.1. The shift space
Endow $\mathscr{D}^{\mathbb{N}}$ with the product topology assuming each factor has a discrete topology. Then, $\mathscr{D}^{\mathbb{N}}$ can be metricized with the metric $\operatorname{dist}$ given by
$\forall\mathbf{a},\mathbf{b}\in\mathscr{D}^{\mathbb{N}}\quad\operatorname{dist}(\mathbf{a},\mathbf{b}):=\begin{cases}0,\text{ if }\mathbf{a}=\mathbf{b},\\ 2^{-\min\{n\in\mathbb{N}:\operatorname{pref}(\mathbf{a};n)\neq\operatorname{pref}(\mathbf{b};n)\}},\text{ if }\mathbf{a}\neq\mathbf{b}.\end{cases}$
@@
Let $\mathbf{a}\in\Omega$ and $\varepsilon>0$ be arbitrary. Take $m\in\mathbb{N}$ such that, for $\psi$ as in Proposition 3.1.iii, $\frac{2}{\psi^{m-1}}<\varepsilon$ and denote by $p_{m}/q_{m}$ the $m$-th convergent of $\Lambda(\mathbf{a})$. If $\mathbf{b}\in\Omega$ is such that $\operatorname{dist}(\mathbf{a},\mathbf{b})<2^{-m}$, then $\operatorname{pref}(\mathbf{a};m)=\operatorname{pref}(\mathbf{b};m)$ and, by Proposition 3.1.v,
@@
In other words, $\Lambda$ is uniformly continuous. Call $\Lambda|_{\mathsf{R}}$ the restriction of $\Lambda$ to $\mathsf{R}$ and let $\overline{\Lambda}$ be the unique continuous extension of $\Lambda|_{\mathsf{R}}$ to $\overline{\mathsf{R}}:=\operatorname{Cl}_{\mathscr{D}^{\mathbb{N}}}(\mathsf{R})$. We may now state a lemma containing the core of Theorem 1.1. We will prove it by constructing algorithmically the sequence $\mathbf{b}$ from $\mathbf{a}$.
Lemma 4.1.
@@
We are now in the position of proving Theorem 1.1.
@@
Pick any $\mathbf{a}=(a_{n})_{n\geq 1}\in\overline{\mathsf{R}}$ and let $(\mathbf{b}^{n})_{n\geq 1}$ be a sequence in $\mathsf{R}$ converging to $\mathbf{a}$. Then, $\overline{\Lambda}(\mathbf{a})\in\overline{\mathfrak{F}}$ because $\overline{\Lambda}$ is continuous and $\overline{\Lambda}(\mathbf{b}^{n})=\Lambda(\mathbf{a})\in\mathfrak{F}$ for all $n\in\mathbb{N}$. Next, we show $\overline{\Lambda}(\mathbf{a})\in\overline{\mathfrak{F}}\setminus\in\mathbb{Q}(i)$ by obtaining infinitely many coprime pairs $(p,q)\in\mathbb{Z}[i]$ satisfying
@@
Pick $N\in\mathbb{N}$ and write $p_{N}=p_{N}(a_{1},\ldots,a_{N})$, $q_{N}=q_{N}(a_{1},\ldots,a_{N})$. For all large $n\in\mathbb{N}$, we have $\operatorname{pref}(\mathbf{b}^{n};N)=\operatorname{pref}(\mathbf{a};N)$ and, by Proposition 3.1.v,
@@
Letting $n\to\infty$ and using the continuity of $\overline{\Lambda}$, we see that $(p,q)=(p_{N},q_{N})$ solves (10). Furthermore, by Proposition 3.1.ii and (6), the pairs $(p_{N},q_{N})$, $N\in\mathbb{N}$, are coprime and different. Therefore, $\overline{\Lambda}(\mathbf{a})\in\overline{\mathfrak{F}}\setminus\mathbb{Q}(i)$.
Now take any $z\in\overline{\mathfrak{F}}\setminus\mathbb{Q}(i)$. Assume that $z=[0;a_{1},a_{2},\ldots]\in\mathfrak{F}$. If $\mathbf{a}\in\mathsf{R}$, then $z=\overline{\Lambda}(\mathbf{a})$. When $\mathbf{a}\in\mathsf{Ir}$, the sequence $\mathbf{b}$ from Lemma 4.1 verifies $\Lambda(\mathbf{a})=\overline{\Lambda}(\mathbf{b})$. When $z\not\in\mathfrak{F}$, we have either $\operatorname{Re}(z)=\tfrac{1}{2}$ or $\operatorname{Im}(z)=\tfrac{1}{2}$. If $\operatorname{Im}(z)=\frac{1}{2}$, then $\overline{z}=\mathsf{mir}_{1}(z)\in\mathfrak{F}$ and there is some $\mathbf{b}=(b_{n})_{n\geq 1}\in\overline{\mathsf{R}}$ such that $\overline{z}=\Lambda(\mathbf{b})$. Thus, by Proposition 3.4, we have $\mathbf{c}:=(\mathsf{mir}_{1}(b_{j}))_{j\geq 1}\in\overline{\mathsf{R}}$ and $\overline{\Lambda}(\mathbf{c})=z$. A similar argument replacing $\mathsf{mir}_{1}$ with $\mathsf{mir}_{2}$ works when $\operatorname{Re}(z)=\frac{1}{2}$. ∎
@@
4.2. The algorithm
@@
$S(a)\colon=\begin{cases}im,\text{ if }a\in\{1+im,-1+im\},\\ m,\text{ if }a\in\{m+i,m-i\}.\end{cases}$
Algorithm 4.3.
@@
    Define the sequence $\mathbf{b}_{N+1}=(b_{N+1}(j))_{j\geq 1}$ as follows:
@@
    If there is no $j_{N}\in\mathbb{N}$ such that (11) holds, define $\mathbf{b}_{N+1}\colon=\mathbf{b}_{N}$. Put $N:=N+1$.
@@
    Take $\mathbf{b}:=\displaystyle\lim_{N\to\infty}\mathbf{b}_{N}$. $\blacksquare$
@@ 4.3. Examples
For the sake of clarity, we illustrate the algorithm with a couple of examples.
@@
so $(-2,1+iB_{1})\in\mathsf{Ir}(2)$ and $[\zeta_{2},\zeta_{1}]=\overline{\mathcal{C}}_{1}(-2)\cap\overline{\mathfrak{F}}_{2}(-2,1+iB_{1})\subseteq\mathfrak{F}_{2}(-2,1+iB_{1})$ (cfr. Figure 2), which implies
@@
As a consequence, $z=[0;a_{1},a_{2},\ldots]_{\mathbb{C}}$ and $\mathbf{a}\in\mathsf{Ir}$, because $(-2,1+iB_{1})\in\mathsf{Ir}(2)$. The algorithm applied to $\mathbf{a}$ gives us the sequence
@@
We may replace the condition $\min\{|B_{n}|:n\in\mathbb{N}\}=3$ by the optimal condition $\min\{|B_{n}|:n\in\mathbb{N}\}=2$ at the expense of some notational complications.
@@
Algorithm 4.3 applied on the extremely irregular sequence

### v3

@@ 3.1. HCFs Cylinders
Naturally, #$\#A$ represents the number of elements in a set $A$. We denote by $\Omega(n)$ (resp. $\mathsf{R}(n)$, $\mathsf{Ir}(n)$, $\mathsf{EI}(n)$) the set of valid (resp. regular, irregular, extremely irregular) words of length $n$. We extend the notions regular, irregular, and extremely irregular to cylinders and prototype sets in an obvious way. It is well known that there are only finitely many prototype sets (see, for example, [16]). Actually, an open prototype set is one of the following sets or one of their right-angled rotations (see Figure 1):
@@
A significant part of our calculations is based on an observation stemming from the definitions of cylinders: if $(a_{n})_{n\geq 1}$ is a sequence in $\mathscr{D}$, then
$\mathcal{C}_{1}(a_{1})=\mathfrak{F}\cap\iota\tau_{a_{1}}[\mathfrak{F}],\quad\mathcal{C}_{2}(a_{1},a_{2})=\mathfrak{F}\cap\iota\tau_{a_{1}}[\mathfrak{F}]\cap\iota\tau_{a_{1}}\iota\tau_{a_{2}}[\mathfrak{F}];$
@@
Since $\iota$ and $\tau_{a}$ can be extended to homeomorphisms from the Riemann sphere onto itself, we may replace $\mathfrak{F}$ and $\mathcal{C}$ with $\mathfrak{F}^{\circ}$ and $\mathcal{C}^{\circ}$, respectively, in (9). It follows from (9) that every factor of a regular word (finite or infinite) is again regular. We also deduce the next proposition.
@@ 3.2. Regular cylinders
We recall two crucial properties of regular cylinders. The first property tells us that every infinite sequence in $\mathscr{D}$ with large terms is regular. The second proposition establishes symmetries of the family of regular cylinders.
@@
and call $\mathsf{Di}_{8}$ the group of isometries they generate. It is clear that $\mathsf{Di}_{8}$ is the dihedral group of order $8$. Proposition 4.5 and Corollary 4.6 in [9] say that any $f\in\mathsf{Di}_{8}$ maps open cylinders onto open cylinders of the same level. These results are proven by explicitly determining the image of an open cylinder under $\mathsf{rot}$ and $\mathsf{mir}_{1}$. In particular, if $\mathsf{mir}_{2}:=\mathsf{rot}^{2}\mathsf{mir}_{1}$, $z\mapsto-\overline{z}$, and $\mathsf{Mir}\subseteq\mathsf{Di}_{8}$ is the subgroup generated by $\mathsf{mir}_{1}$ and $\mathsf{mir}_{2}$, we have the next result.
@@ 3.3. Irregular cylinders
The proof amounts to checking that every option for non-empty $\mathfrak{F}^{\circ}_{n}(\mathbf{a})$, which are not listed in the statement, satisfies
$\mathbf{a}b\in\Omega(n+1)\quad\text{ implies }\quad\mathbf{a}b\in\mathsf{R}(n+1).$
It follows from the stronger statement
@@
For clarity, we rely on pictures to avoid long yet elementary computations.
@@
    When $\mathfrak{F}^{\circ}_{n}(\mathbf{a})=\mathfrak{F}^{\circ}\setminus\overline{\mathbb{D}}(1+i)$ we have $\iota[\overline{\mathfrak{F}}_{n}(\mathbf{a})]=\iota[\overline{\mathfrak{F}}]\setminus\mathbb{D}(1-i)$. We conclude from Figure 4 that $\iota\left[\overline{\mathfrak{F}}_{n}(\mathbf{a})\right]\cap\tau_{b}[\mathfrak{F}]\neq\varnothing$ if and only if $b\neq 1-i$, in which case $\iota\left[\mathfrak{F}^{\circ}_{n}(\mathbf{a})\right]\cap\tau_{b}[\mathfrak{F}^{\circ}]\neq\varnothing$. A similar argument holds when $\mathfrak{F}^{\circ}_{n}(\mathbf{a})=\mathfrak{F}^{\circ}\setminus i^{k}\overline{\mathbb{D}}(1+i)$ for some $k\in\{2,3,4\}$.
@@
    and, by Figure 5, $\iota\left[\overline{\mathfrak{F}}_{n}(\mathbf{a})\right]\cap\tau_{b}\left[\mathfrak{F}\right]\neq\varnothing$ if and only if $\operatorname{Re}(b)\geq 0$. Figure 5 also shows that $\operatorname{Re}(b)\geq 0$ yields $\iota[\mathfrak{F}^{\circ}_{n}(\mathbf{a})]\cap\tau_{b}[\mathfrak{F}^{\circ}]\neq\varnothing$. A similar argument holds when $\mathfrak{F}^{\circ}_{n}(\mathbf{a})=\mathfrak{F}^{\circ}\setminus\overline{\mathbb{D}}(i)$.
@@
    When $\mathfrak{F}^{\circ}_{n}(\mathbf{a})=\mathfrak{F}^{\circ}\setminus(\overline{\mathbb{D}}(i)\cup\overline{\mathbb{D}}(-1))$, by Figure 6 we have $\iota\left[\overline{\mathfrak{F}}_{n}(\mathbf{a})\right]\cap\tau_{b}\left[\mathfrak{F}\right]\neq\varnothing$ if and only if $\operatorname{Re}(b)\geq 0$ and $\operatorname{Im}(b)\geq 0$. For such $b$ we have $\iota[\mathfrak{F}^{\circ}_{n}(\mathbf{a})]\cap\tau_{b}[\mathfrak{F}^{\circ}]\neq\varnothing$.
This shows that if $\mathbf{a}\in\mathsf{R}(n)$ and $\mathfrak{F}^{\circ}_{n}(\mathbf{a})\neq\varnothing$ is not one of the sets listed in the lemma, then $\mathbf{a}b\in\Omega(n+1)$ implies $\mathbf{a}b\in\mathsf{R}(n+1)$. ∎
@@ 4.1. The shift space
Endow $\mathscr{D}^{\mathbb{N}}$ with the product topology assuming each factor has the discrete topology. Then $\mathscr{D}^{\mathbb{N}}$ can be metricized with the metric $\operatorname{dist}$ given by
$\forall\mathbf{a},\mathbf{b}\in\mathscr{D}^{\mathbb{N}}\quad\operatorname{dist}(\mathbf{a},\mathbf{b}):=\begin{cases}0,&\text{ if }\quad\mathbf{a}=\mathbf{b},\\ 2^{-\min\{n\in\mathbb{N}:\operatorname{pref}(\mathbf{a};n)\neq\operatorname{pref}(\mathbf{b};n)\}},&\text{ if }\quad\mathbf{a}\neq\mathbf{b}.\end{cases}$
@@
Let $\mathbf{a}\in\Omega$ and $\varepsilon>0$ be arbitrary. For $\psi$ as in Proposition 3.1.(iii), take $m\in\mathbb{N}$ such that $\frac{2}{\psi^{m-1}}<\varepsilon$ and denote by $p_{m}/q_{m}$ the $m$th convergent of $\Lambda(\mathbf{a})$. If $\mathbf{b}\in\Omega$ is such that $\operatorname{dist}(\mathbf{a},\mathbf{b})<2^{-m}$, then $\operatorname{pref}(\mathbf{a};m)=\operatorname{pref}(\mathbf{b};m)$ and, by Proposition 3.1.(iv),
@@
In other words, $\Lambda$ is uniformly continuous. Call $\Lambda|_{\mathsf{R}}$ the restriction of $\Lambda$ to $\mathsf{R}$ and let $\overline{\Lambda}$ be the unique continuous extension of $\Lambda|_{\mathsf{R}}$ to $\overline{\mathsf{R}}:=\operatorname{Cl}_{\mathscr{D}^{\mathbb{N}}}(\mathsf{R})$. Note that this definition of $\overline{\mathsf{R}}$ is equivalent to the one in the introduction. We now state Proposition 4.1, which contains the core of Theorem 1.1. We prove it by constructing the sequence $\mathbf{b}$ with an iterative process (Procedure 4.3).
Proposition 4.1.
@@
The proof of Proposition 4.1 is in Subsection 7.2. We now show Theorem 1.1.
@@
Pick any $\mathbf{a}=(a_{n})_{n\geq 1}\in\overline{\mathsf{R}}$ and let $(\mathbf{b}^{n})_{n\geq 1}$ be a sequence in $\mathsf{R}$ converging to $\mathbf{a}$. Then, $\overline{\Lambda}(\mathbf{a})\in\overline{\mathfrak{F}}$ because $\overline{\Lambda}$ is continuous and $\overline{\Lambda}(\mathbf{b}^{n})=\Lambda(\mathbf{b}^{n})\in\mathfrak{F}$ for all $n\in\mathbb{N}$. Next, we show $\overline{\Lambda}(\mathbf{a})\in\overline{\mathfrak{F}}\setminus\in\mathbb{Q}(i)$ by obtaining infinitely many coprime pairs $(p,q)\in\mathbb{Z}[i]$ satisfying
@@
Pick $N\in\mathbb{N}$ and write $p_{N}=p_{N}(a_{1},\ldots,a_{N})$, $q_{N}=q_{N}(a_{1},\ldots,a_{N})$. For all large $n\in\mathbb{N}$, we have $\operatorname{pref}(\mathbf{b}^{n};N)=\operatorname{pref}(\mathbf{a};N)$ and, by Proposition 3.1.(iv),
@@
Letting $n\to\infty$ and using the continuity of $\overline{\Lambda}$, we see that $(p,q)=(p_{N},q_{N})$ solves (11). Furthermore, by Proposition 3.1.(ii) and (7), the pairs $(p_{N},q_{N})$, $N\in\mathbb{N}$, are coprime and different. Therefore, $\overline{\Lambda}(\mathbf{a})\in\overline{\mathfrak{F}}\setminus\mathbb{Q}(i)$.
Now consider $z\in\overline{\mathfrak{F}}\setminus\mathbb{Q}(i)$. Assume that $z=[0;a_{1},a_{2},\ldots]\in\mathfrak{F}$. If $\mathbf{a}=(a_{n})_{n\geq 1}\in\mathsf{R}$, then $z=\overline{\Lambda}(\mathbf{a})$. When $\mathbf{a}\in\mathsf{Ir}$, the sequence $\mathbf{b}$ from Proposition 4.1 verifies $\Lambda(\mathbf{a})=\overline{\Lambda}(\mathbf{b})$. When $z\not\in\mathfrak{F}$, we have either $\operatorname{Re}(z)=\tfrac{1}{2}$ or $\operatorname{Im}(z)=\tfrac{1}{2}$. If $\operatorname{Im}(z)=\frac{1}{2}$, then $\overline{z}=\mathsf{mir}_{1}(z)\in\mathfrak{F}$ and there is some $\mathbf{b}=(b_{n})_{n\geq 1}\in\overline{\mathsf{R}}$ such that $\overline{z}=\overline{\Lambda}(\mathbf{b})$. Thus, by Proposition 3.4, we have $\mathbf{c}:=(\mathsf{mir}_{1}(b_{j}))_{j\geq 1}\in\overline{\mathsf{R}}$ and $\overline{\Lambda}(\mathbf{c})=z$. A similar argument replacing $\mathsf{mir}_{1}$ with $\mathsf{mir}_{2}$ works when $\operatorname{Re}(z)=\frac{1}{2}$. ∎
@@
4.2. The construction of $\mathbf{b}$
In this subsection, we provide a procedure that gives the sequence $\mathbf{b}$ of Proposition 4.1.
@@ 4.2. The construction of $\mathbf{b}$
$S(a)\colon=\begin{cases}im,&\text{ if }\quad a\in\{1+im,-1+im\},\\ m,&\text{ if }\quad a\in\{m+i,m-i\}.\end{cases}$
Procedure 4.3.
To any $\mathbf{a}\in\mathsf{Ir}$ we associate $\mathbf{b}\in\mathscr{D}^{\mathbb{N}}$ as in Proposition 4.1.
@@
    Define the sequence $\mathbf{b}_{N+1}=(b_{N+1}(l))_{l\geq 1}$ as follows:
@@
    If there is no $j_{N}\in\mathbb{N}$ such that (12) holds, define $\mathbf{b}_{N+1}\colon=\mathbf{b}_{N}$. Update $N:=N+1$.
@@
    Take $\mathbf{b}:=\displaystyle\lim_{N\to\infty}\mathbf{b}_{N}$.
The reader might wonder whether this process can be performed and, if so, whether the output $\mathbf{b}$ has the desired properties. We show in Section 7 the convergence of $(\mathbf{b}_{N})_{N\geq 1}$, $\mathbf{b}\in\overline{\mathsf{R}}$, and $\overline{\Lambda}(\mathbf{b})=\Lambda(\mathbf{a})$.
@@ 4.3. Examples
For clarity, we illustrate the previous procedure with a couple of examples.
@@
so $(-2,1+iB_{1})\in\mathsf{Ir}(2)$ and $[\zeta_{2},\zeta_{1}]=\overline{\mathcal{C}}_{1}(-2)\cap\overline{\mathfrak{F}}_{2}(-2,1+iB_{1})\subseteq\mathfrak{F}_{2}(-2,1+iB_{1})$ (see Figure 2), which implies
@@
As a consequence, $z=[0;a_{1},a_{2},\ldots]_{\mathbb{C}}$ and $\mathbf{a}\in\mathsf{Ir}$, because $(-2,1+iB_{1})\in\mathsf{Ir}(2)$. Procedure 4.3 applied to $\mathbf{a}$ gives us the sequence
@@
We may replace the condition $\min\{|B_{n}|:n\in\mathbb{N}\}=3$ by the optimal condition $\min\{|B_{n}|:n\in\mathbb{N}\}=2$ at the expense of notational complications.
@@
Procedure 4.3 applied on the extremely irregular sequence